# The purpose of an estimator is what it does: Misspecification, estimands, and over-identification

### v3

@@ 2. What Do We Mean by “Misspecified?”
If the econometric model is correct, least squares regression of $Y_{i}$ on $W_{i}$ recovers $\theta,$ so we can define the statistical parameter as $\theta^{*}(P)=\frac{\operatorname{Cov}_{P}(W_{i},Y_{i})}{\operatorname{Var}_{P}(Y_{i})}.$ $\blacksquare$
@@ 2.1. An aside on the many meanings of models
may model linear, homogeneous causal effects $Y_{i}(w)=\alpha+\beta w+\epsilon_{i}$; or it may simply define $\beta$ as a statistical functional equal to the population OLS coefficient, leaving the joint distribution of $(Y,W)$ unrestricted. Under our terminology, the two cases correspond to distinct econometric models. [^11] The former describes a model from homogeneous and linear potential outcomes $\mathcal{M}=\left\{{(\beta,P):\times_{i=1}^{n}(Y_{i}(W),W_{i})\sim P,Y_{i}(w_{1})-Y_{i}(w_{2})=\beta(w_{1}-w_{2})}\right\}.$ The latter defines a model in which $\mathcal{M}=\left\{{\left({\frac{\operatorname{Cov}_{P}(W,Y)}{\operatorname{Var}_{P}(Y)},P}\right):\times_{i=1}^{n}(Y_{i},W_{i})\sim P,\text{ $P$ has second moments}}\right\}.$
@@ 3. Choice of Estimator and Estimand
Moment condition models are usually estimated by the generalized method of moments (GMM). For $g_{n}(X^{n},\theta)=\frac{1}{n}\sum g(X_{i},\psi)$ the sample average moments and $\hat{\Omega}=\Omega_{n}(X^{n})$ a data-dependent weighting matrix, GMM takes $\hat{\theta}_{\Omega}=\vartheta\left(\hat{\psi}_{\Omega}\right)$ for
@@
As previously discussed, under misspecification and provided regularity conditions hold, the GMM estimator $\hat{\psi}_{\Omega}$ converges to the value which minimizes the $\Omega_{P}$-weighted deviation of the moment conditions from zero. In contrast, $\text{plim}(\hat{\psi}_{\text{MLE}},P)=\operatorname*{arg\,min}_{\psi\in\Psi}\mathbb{E}_{P}[\log p_{\psi}(X^{n})],$ which minimizes the Kullback–Leibler divergence between the distribution of the data and the model-implied distribution (White, 1982). $\blacksquare$
@@
One reason the “optimal” weighting matrix may not be used in some cases is that in settings where the sample size is small relative to the number of moments, it is known that estimation of the optimal weighting matrix can lead to bias in both the estimate and the standard errors (see e.g. Altonji and Segal, 1996; Windmeijer, 2005). One suggestion in such cases is to use the inverse of the diagonal alone, which Altonji and Segal (1996) found yielded “a substantial bias reduction” their problem. [^16] Consistent with this intepretation, Abebe et al. (2021) write that “In line with the recent literature, we use this simple weighting matrix instead of the theoretically optimal weighting matrix, which may suffer from small sample bias (Altonji and Segal, 1996).” However, together with optimal weighting this still covers only about half of the over-identified papers in our GMM sample.
@@ 4. Inference on Estimands
When the model is well-specified $g_{n}(\psi_{\Omega}^{*})$ has mean zero, so the second term in (3) is negligible, $\sqrt{n}(\hat{\Gamma}^{\prime}\hat{\Omega}-\Gamma^{\prime}\Omega)g_{n}(\psi_{\Omega}^{*})=o_{P}(1)$, and the estimator is asymptotically normal with the conventional GMM asymptotic variance:
@@ 5.1. J-Statistics, not J-Tests
Relatively few economics papers report J-statistics. In our GMM sample, for instance, only 3 out of the 22 over-identified papers report J-statistics, either in the paper or the online appendix. One reason may be that few economists expect their models to hold exactly. Under that view, a non-rejection signals insufficient power of the J-test, not correct specification. This suggests a focus on “economic” rather than statistical measures of model fit. Consistent with this, 19 of the 22 over-identified papers in our GMM sample report some version of an informal identification check, often by plotting the model-implied values of a set of moments against either (i) the data moments that were used to estimate the model or (ii) a “held out” set of moments not used in estimation, which may be interpreted as estimation moments assigned zero weight by $\hat{\Omega}$. Both of these approaches convey information about the extent to which the model’s over-identifying restrictions appear to hold in the sample, but lack the formal statistical interpretation of the J-statistic.
Such informal checks, while potentially useful, are an imperfect substitute for J-statistics. To highlight this point, in this section we show that J-statistics are useful for much more than J-tests. As discussed in the introduction, the fact that researchers may choose different weighting matrices due to misspecification concern introduces an additional researcher degree of freedom. We show below that (under local misspecification) J-statistics capture, in a precise sense, the range of estimates, t-statistics, and confidence sets a researcher may obtain by varying their choice of weight matrix. This is valuble information for readers, and we encourage wider reporting of J-statistics in applications of GMM.
@@
$\hat{\sigma}_{\Omega}^{2}=\frac{1}{n}\hat{h}^{\prime}\left(\hat{\Gamma}_{\Omega}^{\prime}\Omega\hat{\Gamma}_{\Omega}\right)^{-1}\hat{\Gamma}_{\Omega}^{\prime}\Omega\hat{\Sigma}_{\Omega}\Omega\hat{\Gamma}_{\Omega}\left(\hat{\Gamma}_{\Omega}^{\prime}\Omega\hat{\Gamma}_{\Omega}\right)^{-1}\hat{h}\quad\hat{h}=\frac{\partial{\varphi(\hat{\psi}_{\Omega})}}{\partial{\psi}}$
@@ A.1. Proofs for Normal Limit Problem
$\sup_{v_{\Omega}^{\prime}\Sigma v_{\Omega}\leq\tau^{2}\sigma_{\Sigma^{-1}}^{2}}v_{\Omega}^{\prime}Z=\tau\sigma_{\Sigma^{-1}}\lVert\Sigma^{-1/2}Z\rVert=\tau\sigma_{\Sigma^{-1}}\sqrt{J}$
@@
The confidence set $C_{\Omega}(c)$ has radius $c\sqrt{1+\tau^{2}}\sigma_{\Sigma^{-1}}<\tau\sqrt{J}\sigma_{\Sigma^{-1}}$ for sufficiently large $\tau^{2}$. Thus, for sufficiently large $\tau^{2}$, the confidence set $C_{\Omega}(c)$ would fail to include $\hat{\theta}_{\Omega}$. On the other hand, for such a $\tau$, there is another estimator that is on the other side of $\hat{\theta}_{\Sigma^{-1}}$: $\hat{\theta}_{\Omega^{\prime}}=\hat{\theta}_{\Sigma^{-1}}-\sigma_{\Sigma^{-1}}\tau\sqrt{J}$. Hence $C_{\Omega}(c)\cap C_{\Omega^{\prime}}(c)=\varnothing$ since neither extends to $\hat{\theta}_{\Sigma^{-1}}$. For $c=\sqrt{J}$,
@@
This is immediate from Corollary A.4. In particular, note that by Corollary A.4, $t_{\omega}(\hat{\theta}_{\Sigma^{-1}})\leq\sqrt{J}$ for all $\Omega,$ while for any $\theta\neq\hat{\theta}_{\Sigma^{-1}})$ there exists $\Omega$ with $t_{\omega}(\hat{\theta}_{\Sigma^{-1}})>\sqrt{J}$. ∎
@@
$\Omega=\Gamma\Gamma^{\prime}+a(uw^{\prime}+wu^{\prime})+ww^{\prime}+P_{V}$
@@
Hence $\Omega$ is positive definite:
$t^{\prime}\Omega t=t_{1}(\Gamma^{\prime}\Gamma)^{2}t_{1}+c^{2}\lVert w\rVert^{2}+\lVert t_{2}\rVert^{2}>0.$
@@
$\Omega\Gamma(\Gamma^{\prime}\Omega\Gamma)^{-1}=\Gamma(\Gamma^{\prime}\Gamma)^{-1}+awh^{\prime}(\Gamma^{\prime}\Gamma)^{-2}.$
Choose $a=\frac{1}{h^{\prime}(\Gamma^{\prime}\Gamma)^{-2}h}>0$ so that
$\Omega\Gamma(\Gamma^{\prime}\Omega\Gamma)^{-1}h=\Gamma(\Gamma^{\prime}\Gamma)^{-1}h+w=u+w=q$
@@ Appendix B Proofs for Results in Main Text
$\sqrt{n}(\hat{\theta}_{\Omega}-\vartheta(\psi))-\hat{\theta}^{*}_{\Omega}\overset{p}{\longrightarrow}0,~~\sqrt{n}\hat{\sigma}_{\Omega}\overset{p}{\longrightarrow}\sigma_{\Omega}=h^{\prime}\left(\Gamma^{\prime}\Omega\Gamma\right)^{-1}\Gamma^{\prime}\Omega\Sigma\Omega\Gamma\left(\Gamma^{\prime}\Omega\Gamma\right)^{-1}h,$ (9)
@@
$\lim_{\kappa\to 0}\operatorname{P}\left\{N\left(\eta,\Sigma\right)\in\mathcal{G}(\kappa,\zeta)\right\}=1,$
@@
$\lim_{\kappa\to\infty}\limsup_{n\to\infty}\operatorname{P}\left\{\sup_{\Omega\in\mathcal{W}_{\kappa}}\left|t_{\Omega}\left(\hat{\theta}_{\Sigma^{-1}}+\frac{1}{\sqrt{n}}\delta\right)\right|>\frac{\tau}{\sqrt{1+\tau^{2}}}J+\frac{|\delta|}{\sqrt{1+\tau^{2}}\sigma_{\Sigma^{-1}}}-\nu^{\prime}\right\}=1.$
@@
$\lim_{\nu\to 0}\sup_{\tau>0}\liminf_{n\to\infty}\operatorname{P}\left\{J+\nu<\frac{\tau}{\sqrt{1+\tau^{2}}}\sqrt{J}+\frac{|\delta|}{\sqrt{1+\tau^{2}}\sigma_{\Sigma^{-1}}}\right\}=1,$

### v4

@@ 2. What Do We Mean by “Misspecified?”
If the econometric model is correct, least squares regression of $Y_{i}$ on $W_{i}$ recovers $\theta,$ so we can define the statistical parameter as $\theta^{*}(P)=\frac{\operatorname{Cov}_{P}(W_{i},Y_{i})}{\operatorname{Var}_{P}(W_{i})}.$ $\blacksquare$
@@ 2.1. An aside on the many meanings of models
may model linear, homogeneous causal effects $Y_{i}(w)=\alpha+\beta w+\epsilon_{i}$; or it may simply define $\beta$ as a statistical functional equal to the population OLS coefficient, leaving the joint distribution of $(Y,W)$ unrestricted. Under our terminology, the two cases correspond to distinct econometric models. [^11] The former describes a model from homogeneous and linear potential outcomes $\mathcal{M}=\left\{{(\beta,P):\bigtimes_{i=1}^{n}(Y_{i}(W),W_{i})\sim P,Y_{i}(w_{1})-Y_{i}(w_{2})=\beta(w_{1}-w_{2})}\right\}.$ The latter defines a model in which $\mathcal{M}=\left\{{\left({\frac{\operatorname{Cov}_{P}(W,Y)}{\operatorname{Var}_{P}(W)},P}\right):\bigtimes_{i=1}^{n}(Y_{i},W_{i})\sim P,\text{ $P$ has second moments}}\right\}.$
@@ 3. Choice of Estimator and Estimand
Moment condition models are usually estimated by the generalized method of moments(GMM). For $g_{n}(X^{n},\psi)=\frac{1}{n}\sum g(X_{i},\psi)$ the sample average moments and $\hat{\Omega}=\Omega_{n}(X^{n})$ a data-dependent weighting matrix, GMM takes $\hat{\theta}_{\Omega}=\vartheta\left(\hat{\psi}_{\Omega}\right)$ for
@@
As previously discussed, under misspecification and provided regularity conditions hold, the GMM estimator $\hat{\psi}_{\Omega}$ converges to the value which minimizes the $\Omega_{P}$-weighted deviation of the moment conditions from zero. In contrast, $\text{plim}(\hat{\psi}_{\text{MLE}},P)=\operatorname*{arg\,max}_{\psi\in\Psi}\mathbb{E}_{P}[\log p_{\psi}(X^{n})],$ which minimizes the Kullback–Leibler divergence between the distribution of the data and the model-implied distribution (White, 1982). $\blacksquare$
@@
One reason the “optimal” weighting matrix may not be used in some cases is that in settings where the sample size is small relative to the number of moments, it is known that estimation of the optimal weighting matrix can lead to bias in both the estimate and the standard errors (see e.g. Altonji and Segal, 1996; Windmeijer, 2005). One suggestion in such cases is to use the inverse of the diagonal alone, which Altonji and Segal (1996) found yielded “a substantial bias reduction” their problem. [^16] Consistent with this intepretation, Abebe et al. (2021a) write that “In line with the recent literature, we use this simple weighting matrix instead of the theoretically optimal weighting matrix, which may suffer from small sample bias (Altonji and Segal, 1996).” However, together with optimal weighting this still covers only about half of the over-identified papers in our GMM sample.
@@ 4. Inference on Estimands
When the model is well-specified $g_{n}(\psi_{\Omega}^{*})$ has mean zero, so the second term in (3) is negligible, $\sqrt{n}(\hat{\Gamma}^{\prime}\hat{\Omega}-\Gamma^{\prime}\Omega)g_{n}(\psi_{\Omega}^{*})=o_{P}(1)$, and the estimator is asymptoticaly normal with the conventional GMM asymptotic variance:
@@ 5.1. J-Statistics, not J-Tests
Relatively few economics papers report J-statistics. In our GMM sample, for instance, only 3 out of the 22 over-identified papers report J-statistics, either in the paper or the online appendix. [^24] One additional paper, Mueller et al. (2021) reports the value of the two-step SMM objective function at its minimum. Because they use the optimal weighting matrix (estimated by bootstrap), this objective value is equal to the J-statistic absent simulation error. One reason may be that few economists expect their models to hold exactly. Under that view, a non-rejection signals insufficient power of the J-test, not correct specification. This suggests a focus on “economic” rather than statistical measures of model fit. Consistent with this, 19 of the 22 over-identified papers in our GMM sample report some version of an informal identification check, often by plotting the model-implied values of a set of moments against either (i) the data moments that were used to estimate the model or (ii) a “held out” set of moments not used in estimation, which may be interpreted as estimation moments assigned zero weight by $\hat{\Omega}$. Both of these approaches convey information about the extent to which the model’s over-identifying restrictions appear to hold in the sample, but lack the formal statistical interpretation of the J-statistic.
Such informal checks, while potentially useful, are an imperfect substitute for J-statistics. To highlight this point, in this section we show that J-statistics are useful for much more than J-tests. As discussed in the introduction, the fact that researchers may choose different weighting matrices due to misspecification concern introduces an additional researcher degree of freedom. We show below that (under local misspecification) J-statistics capture, in a precise sense, the range of estimates, t-statistics, and confidence sets a researcher may obtain by varying their choice of weight matrix. This is valuable information for readers, and we encourage wider reporting of J-statistics in applications of GMM.
To better understand the behavior of J statistics in practice, we attempted to calculate J statistics for the papers in our GMM sample which did not already report them. In this exercise we were limited to those papers with accessible data, and where the optimally-weighted GMM objective was sufficiently well-behaved that we were confident in the convergence of our numerical optimization routines. Subject to these restrictions, we were able to compute J-statistics for an additional three papers beyond those which initially reported them. Table 1 collects the J-statistics for these papers, along with the original three. Strikingly, for the 6 papers in our sample where we were able to obtain J statistics, in only one case is the model rejected at the 5% level (although three papers have specifications rejected at the 10% level). Also striking, while the number of papers is limited, there is not obvious selection in which papers report J-statistic. Indeed, the papers for which we computed J-statistics have, on the whole, larger p-values (and thus, in that sense, less evidence of misspecification) than the papers which originally reported J-statistics.
Table 1. $X$
Paper Model J-statistic dF p-value Reported in the paper?
Table 4, Column 1 3.05 1 0.08
Abebe et al. (2021b) Table 4, Column 2 4.67 3 0.20 No
Table 4, Column 4 4.28 2 0.12
Bourreau et al. (2021) Table 4, Column 2 50.88 7 0.00 J-statistic and dF
Mueller et al. (2021) Table 5 0.462 3 0.93 No
García-Jimeno et al. (2022) Table 2, Column 2 2.16 5 0.83 J-statistic and p-value
Table 6, Column 7 0.361 1 0.55
Karahan et al. (2024) Table 7, Column 4 5.813 2 0.05 p-value
Table 7, Column 5 0.327 1 0.57
Kelley et al. (2024) Table 7 0.005 1 0.94 No
  • Notes: This table reports J-statistics and associated degrees of freedom and p-values for six papers in our “GMM sample.” See the text for inclusion criteria.
@@
$\hat{\sigma}_{\Omega}^{2}=\frac{1}{n}\hat{h}^{\prime}\left(\hat{\Gamma}_{\Omega}^{\prime}\Omega\hat{\Gamma}_{\Omega}\right)^{-1}\hat{\Gamma}_{\Omega}^{\prime}\Omega\hat{\Sigma}_{\Omega}\Omega\hat{\Gamma}_{\Omega}\left(\hat{\Gamma}_{\Omega}^{\prime}\Omega\hat{\Gamma}_{\Omega}\right)^{-1}\hat{h}\quad\hat{h}=\frac{\partial{\vartheta(\hat{\psi}_{\Omega})}}{\partial{\psi}}$
@@
Absent restrictions on the weighting matrix the J-statistic threshold necessary for a result to be “weight-hacking” proof at the 5% level, $\sqrt{J}\leq 1.96$ does not scale with the degree of over-identification, and so can be much more stringent than conventional J-test critical values. In Table 1, for instance, four of ten specifications, and 3 of 6 papers, have $\sqrt{J}>1.96$, despite only one specification being rejected at the 5% level by the $J$-test.
@@
Returning to Table 1, of the 10 specifications, 4 have J-statistics larger than $1.96^{2}=3.84$; for these specifications, it is possible to choose weight matrices to reject any point null hypothesis. More quantitatively, using 5.2, we can interpret conventional confidence sets around the efficient estimator $\hat{\theta}_{\Sigma^{-1}}$ as a set of GMM estimates at a given standard error inflation $\sqrt{1+\tau^{2}}-1$ for $\tau=1.96/\sqrt{J}$. The interquartile range of this implied standard error inflation is 35% to 230%.
@@ A.1. Proofs for Normal Limit Problem
$\sup_{v_{\Omega}^{\prime}\Sigma_{Z}^{*}v_{\Omega}\leq\tau^{2}\sigma_{\Sigma^{-1}}^{2}}v_{\Omega}^{\prime}Z=\tau\sigma_{\Sigma^{-1}}\lVert{\Sigma_{Z}^{*}}^{-1/2}Z\rVert=\tau\sigma_{\Sigma^{-1}}\sqrt{J}$
@@
The confidence set $C_{\Omega}(c)$ has radius $c\sqrt{1+\tau^{2}}\sigma_{\Sigma^{-1}}<\tau\sqrt{J}\sigma_{\Sigma^{-1}}$ for sufficiently large $\tau^{2}$. Thus, for sufficiently large $\tau^{2}$, the confidence set $C_{\Omega}(c)$ would fail to include $\hat{\theta}_{\Sigma^{-1}}$. On the other hand, for such a $\tau$, there is another estimator that is on the other side of $\hat{\theta}_{\Sigma^{-1}}$: $\hat{\theta}_{\Omega^{\prime}}=\hat{\theta}_{\Sigma^{-1}}-\sigma_{\Sigma^{-1}}\tau\sqrt{J}$. Hence $C_{\Omega}(c)\cap C_{\Omega^{\prime}}(c)=\varnothing$ since neither extends to $\hat{\theta}_{\Sigma^{-1}}$. For $c=\sqrt{J}$,
@@
This is immediate from Corollary A.4. In particular, note that by Corollary A.4, $t_{\Omega}(\hat{\theta}_{\Sigma^{-1}})\leq\sqrt{J}$ for all $\Omega,$ while for any $\theta$ there exists $\Omega$ with $t_{\Omega}(\theta)\geq\sqrt{J}$. ∎
@@
$\Omega=\Gamma\Gamma^{\prime}+a(uw^{\prime}+wu^{\prime})+aww^{\prime}+P_{V}$
@@
Choose $a=\frac{1}{h^{\prime}(\Gamma^{\prime}\Gamma)^{-2}h}>0$. Observe that $\Omega$ is positive definite:
$t^{\prime}\Omega t$ $=t_{1}(\Gamma^{\prime}\Gamma)^{2}t_{1}+ac^{2}\lVert w\rVert^{4}+\lVert t_{2}\rVert^{2}+2ac\lVert w\rVert^{2}t_{1}^{\prime}h$
$\geq a(t_{1}^{\prime}h)^{2}+ac^{2}\lVert w\rVert^{4}+2ac\lVert w\rVert^{2}t_{1}^{\prime}h+\lVert t_{2}\rVert^{2}$ (By Cauchy–Schwarz, $(t_{1}^{\prime}h)^{2}\leq(t_{1}^{\prime}(\Gamma^{\prime}\Gamma)^{2}t_{1})(h^{\prime}(\Gamma^{\prime}\Gamma)^{-2}h)$ )
$=a(t_{1}^{\prime}h+c\lVert w\rVert^{2})^{2}+\lVert t_{2}\rVert^{2}\geq 0.$
@@
$\Omega\Gamma(\Gamma^{\prime}\Omega\Gamma)^{-1}=\Gamma(\Gamma^{\prime}\Gamma)^{-1}+awh^{\prime}(\Gamma^{\prime}\Gamma)^{-2}\implies\Omega\Gamma(\Gamma^{\prime}\Omega\Gamma)^{-1}h=\Gamma(\Gamma^{\prime}\Gamma)^{-1}h+w=u+w=q$
@@ Appendix B Proofs for Results in Main Text
$\sqrt{n}(\hat{\theta}_{\Omega}-\vartheta(\psi))-\hat{\theta}^{*}_{\Omega}\overset{p}{\longrightarrow}0,~~\sqrt{n}\hat{\sigma}_{\Omega}\overset{p}{\longrightarrow}\sigma_{\Omega}=\left({h^{\prime}\left(\Gamma^{\prime}\Omega\Gamma\right)^{-1}\Gamma^{\prime}\Omega\Sigma\Omega\Gamma\left(\Gamma^{\prime}\Omega\Gamma\right)^{-1}h}\right)^{1/2},$ (9)
@@
$\lim_{\kappa\to\infty}\operatorname{P}\left\{N\left(\eta,\Sigma\right)\in\mathcal{G}(\kappa,\zeta)\right\}=1,$
@@
$\lim_{\kappa\to\infty}\limsup_{n\to\infty}\operatorname{P}\left\{\sup_{\Omega\in\mathcal{W}_{\kappa}}\left|t_{\Omega}\left(\hat{\theta}_{\Sigma^{-1}}+\frac{1}{\sqrt{n}}\delta\right)\right|>\frac{\tau}{\sqrt{1+\tau^{2}}}\sqrt{J}+\frac{|\delta|}{\sqrt{1+\tau^{2}}\sigma_{\Sigma^{-1}}}-\nu^{\prime}\right\}=1.$
@@
$\lim_{\nu\to 0}\sup_{\tau>0}\liminf_{n\to\infty}\operatorname{P}\left\{\sqrt{J}+\nu<\frac{\tau}{\sqrt{1+\tau^{2}}}\sqrt{J}+\frac{|\delta|}{\sqrt{1+\tau^{2}}\sigma_{\Sigma^{-1}}}\right\}=1,$